Considering old nos.like 1,3,5,9,11,13, 15 to take any 5 number from this total should be 30 you can reapet the no.also
1
Expert's answer
2014-06-20T12:35:50-0400
If only 5 numbers are used with assumption that we can apply arithmetical operations, then I can suggest this example: 30=11+15/3+13+1. There are solutions when only 4 (not 5) numbers are used (30=15+13+5-3=15+11+3+1=15+11+5-1), when 6 numbers are used (30=1+3+5+9+13-11) or when we can apply numbers several times (30=15+11+3-1+5-3).